OpenStudy (anonymous):

What is the equation of the parabola with vertex (0,0) and directrix y=2?

OpenStudy (anonymous):

A)y=-1/8x^2 b)y=1/2x^2 c)y=1/8x^2 d)y=-1/2x^2

OpenStudy (anonymous):

@IMStuck help?

OpenStudy (imstuck):

Ok, I'm here to help you with this!

OpenStudy (imstuck):

It's an upside down parabola in the form\[x ^{2}=-4py\]

OpenStudy (anonymous):

would it be answer A then?

OpenStudy (imstuck):

It looks like this: I always draw them out, even after all this time, because looking at it on paper is so much easier! I'm not sure of thee answer yet..I need to still find it!

OpenStudy (imstuck):



OpenStudy (imstuck):

The focus is always the same distance from the vertex as the directrix is. So if the directrix is at y = 2, then the focus is down two from the vertex at (0, -2). This is your p value...-2. Fit it into the equation now:

OpenStudy (imstuck):

\[x ^{2}=-4(-2)y\]which comes to\[x ^{2}=8y\]now divide both sides by 8 and get\[\frac{ 1 }{8 }x ^{2}=y\]That's it!

OpenStudy (imstuck):

It is A. I left off the - sign. Sorry. Good job! Yes it is A.

OpenStudy (anonymous):

Thank you!! @IMStuck you're awesome

OpenStudy (imstuck):

awww....go on......and on and on... ; )